# GiBy: A Giant-Step Baby-Step Classifier For Anomaly Detection In Industrial Control Systems

### v1

@@ VI-A Detector limitations
Our core detection solution (see Section III) is limited to its nn-actuators, for any given sensor. We mapped a smaller state-space of actuator states to the sensor reading. As a result, we forsook the determination of a larger number of tighter individual bounds that takes exponential time for the convenience of a practically smaller solution seek space. However, our sensor measurements are taken directly from the normal operations dataset. As a result, the overall determined [LB, UB] bounds remained unchanged, and hence does not change the overall safety limits determined. For example, consider eight dependent actuators influencing a sensor reading. When the actuation sequence for each actuator is on, off, then . When and number of actuators is 8, there are a total of actuation states. This requires determination of bounds. Additionally, let us assume that only 3 out of the 8 dependent actuators are nearest neighbours, w.r.t. sensor . Using GiBy, we would have only considered nearest neighbour () actuation states, and determined . As a result, there was no change in the overall state sensor bounds (safety limits) determined, although some detection resolution was lost by considering only the nearest neighbours.
@@
Another limitation expected from the anomaly detector is a sensitivity to out of bound measurements, resulting in warnings being issued when no attack is present. This limitation could occur even in the system normal (initial) operating phase. The normal dataset is expected to cover its everyday boundaries of operation. However, in some cases the normal dataset may not cover all boundaries of normal operations, and these normal boundaries may be surpassed during future plant operations. Any logged warning observed, that is deemed as normal, will have to be reviewed and trained, to avoid those warnings from re-appearing. However, this will stabilize once sufficient normal (training) data indicative of its range of operations is available. This is a training dataset limitation that affects our model. A final limitation is that it may not detect all invalid actuation states. It is limited to the resolution of the actuator states seen for the nn-actuators.
VI-B Detection in an early prototype system design
@@
VI-C GiBy Detection capabilities
@@
In the threat model (Section II-D), the attacks are classified into single-stage single/multiple component attacks and multi-stage single/multiple component attacks. The component considered is a sensor or an actuator. The goal of this classification is to build a quantifiable measure (score) on the quality of the detector, w.r.t. its ability to detect these types of attacks. Without loss of generality, the classification may be merged into single and multiple sensor/actuator attack(s) on a system.
@@
Our core detector (see Section III) can detect single sensor or actuator manipulations, when it breaks the safety bounds of the relation between sensor and its actuation state. When both sensor and corresponding actuator(s) are spoofed, it is no longer able to individually detect that attack, but it depends on whether any of the sensor readings downstream222Downstream bounds are also captured by GiBy, since it is computed for every sensor with nn-neighbour actuator(s)., as a result, went out of bounds. For example, in Fig. 3, consider an attack where P101 is spoofed to ON and FIT201 reading to , despite the pump being OFF and no water flowing. Assume the controller logic is to turn on the first HCl dosing pump (P201), when the FIT201 flow rate is between and to OFF otherwise. The attack causes the dosing pump to release HCl even when no water is pumped333 The success of this attack depends on whether the attacker can compromise the controller PLC and manipulate its measurements.. However, due to pH sensor AIT201, this attack is detected. Further, if we assume that AIT201 is also spoofed to show normal range of measurements — a redundant downstream sensor AIT501 incorporated in is able to detect the abnormality in pH. However, if AIT501 is also spoofed, it may not be able to detect the attack from the measurements made. As a side effect, changing the pH might possibly change its conductivity sensor reading, permitting the attack to be detected. It would be difficult to conclude without further experimentation.
In some scenarios, it may be sufficient to detect the anomaly at a later stage, and in others, it requires to be detected in the same stage. For example, if pump P101 is running above its safe rpm, it might cause the adjacent (following) pipe to burst, affecting the plant and personnel. Where such a concern exists, the design has to include a pressure sensor. When there is a risk the pressure sensor reading is spoofed, a physical pressure safety (release) valve will have to be incorporated into the design. An anomaly detector’s ability to detect single/multiple attacks is highly dependent on intrinsic design relationships and system dependencies. This tight coupling makes generalized interpretations on detection capability (by only considering the detection model) less meaningful.
VI-D Non-neighbor PLC imposed actuation
@@
The stage in Fig. 3 involves chemical dosing. Dosing tanks are required to be in the order NaCl, HCl and NaOCl (see Section VI-B for reason). AIT201 in the SWaT dataset measures conductivity. According to the PLC control logic sheet, the conductivity sensor AIT503 (downstream sensor in ) read by the PLC instructs P201 to be turned off, when this sensor value hits a preset threshold. Turning off the salt pump, in-turn regulates the conductivity sensor reading at AIT503. Hence, P201 is additionally considered as a nn-actuator (though not a neighbour) for the sensor AIT503; to capture this relationship. It is added to train the sensor (using GiBy). In general, the PLC logic sheet is examined to include such ‘additional’ nn-actuators, when there is a sensor dependency. Not including this relationship does not change the overall detection bounds captured in training. However, it means that some detection resolution is lost because we are no longer directly capturing the relationship between AIT503 and pump P201.
VI-E Explainablity
@@
The more sophisticated AI/ML models using neural network and deep learning are inherently black box models. They are used to detect anomalies but requires additional support from eXplaianable AI (XAI) models to explain the decision made by the AI detector. However, both the ML and XAI models have individual accuracy limitations and is compounded when used in combination. It also takes longer to train and test these models (see Section V). On the contrary, explainability with the proposed GiBy detector is straightforward. Note that we employed a switchboard in Section III-C. Traceability is offered via the one-to-one and onto map, using the switchboard — also in the inverse direction, from sensor reading (under testing) to its [LB, UB] bounds (determined in the training phase). Our training and testing are carried out per sensor, for all concerned sensors and nn-actuator(s). When an anomaly is detected, this immediately tells us what sensor readings are causing anomalies. Since the sensor is known, we can look at its current and previous actuation states, from the state log. GiBy raises an anomaly warning when i.) the sensor reading (under testing) for the nn-actuator state is out of bound. Our detector is able to provide an explanation remark, whether it was because the sensor reading was or . ii.) when the normal training dataset did not have the nn-state actuator state for the sensor, in it. This may happen for two reasons a.) the actuation state is invalid or b.) that state is yet to be encountered in the plants’ normal operation and hence not yet seen in the normal training dataset. Another advantage of GiBy, is that it does not produce false positives except when the normal dataset is not wide enough to capture rare non-everyday spikes, anomalies or boundary excursions. Its detection boundaries are clear and limited to what it has seen in normal dataset training. Our detector system is designed to offer a level of explainability that is expected to equip the operator to deal with most contingencies. The capability of our system to provide the correct sensor location and response to faulty components/attacks subsequently reduces the plant downtime.
VI-F Usability in other sectors
@@
Anomaly detection and mitigation in some sectors (such as power), in certain scenarios, require an actuation within 10-15 milliseconds from event occurrence, to ensure safety. The GiBy algorithm presented in Section III is universally transferable and GiBy-core testing may be parallelized. This is because it produces a series of independent [LB, UB] boundaries. It may be pre-programmed into the control logic of an Intelligent Electronic Device (IED), typically used in a power substation. It may be used to trip the circuit breaker in near real-time (when sensor readings are out of safety bounds), or to inform an operator regarding the anomalous behaviour. The computational cost of GiBy-core tests (see Section III-E) are to see whether the sensor value is in between two bounds [LB, UB]; taking up to two comparisons. The main costs of the extended algorithm tests (see Section IV) are the binary search and sliding window multiplications in each time step. Its time complexity may be reduced to that of binary search on the pre-sorted linearized group (that takes logarithmic time), and a maximum of one multiplication and division: for each time step. Instead of repeatedly multiplying every sensor derived value in the sliding window, the newest time step sensor derived value may be multiplied into the existing product, and the sensor derived value at the tail of the window is divided out. As a result, it might also find (outlier detection) applications in resource constrained devices, edge computing devices and legacy systems with lower computational power.
VII Related Work
@@
While there are solutions to detect packet anomalies on the OT network [25] and memory anomalies on computation devices [26, 27], amidst other detection vectors, the focus of this work is on process anomaly detection [28, 29]. Several approaches are available for design-centric and data-centric anomaly detection. In a design-centric approach, for example, Adepu and Mathur [30] captured the design using state condition graphs. It was used to capture the conditions on actuation w.r.t. sensor readings, expressed using Boolean conditions, and represented in a graphical format. The work in Yoong et al. [31] used axiomatic design methodology for systems, where it iteratively decomposes a CPS design to sets of dependent components and transformed into invariants. In Merwa et al. [32], an association rule mining technique was used to generate attack and invariant rules. In a data-centric approach, for example, there exist solutions to distinguish outliers, such as ECOD in Li et al. [10]. It computed a univariate empirical cumulative distribution function for each dimension separately. To measure the chance of a data point being anomalous, they computed its tail probability across all dimensions. Another outlier detector called deep Support Vector Data Description (Deep-SVDD) proposed in Ruff et al. [11], trained a neural network while minimizing the volume of a hypersphere that enclosed the network representations of the data. In WaXAI, Mathuros et al. [12] employed ECOD and Deep-SVDD in the context of SWaT anomaly detection. In addition, they also employed explaianable AI [14] models, namely, kernel SHAP [33], LIME [34], ALE [35] and IG [36], to provide explanation for the anomaly detected. Further, a 20% improvement in attribution of attack root cause over SHAP was achieved by using a Factorization Machines based approach by Avdalovic et al. [37].
In this work we provided attribution (explainability) linked directly to the sensor. Explainability was provided for each sensor when out of bounds, to top up the recommendations in Fung et al [38]. The GiBy classifier (see Section III) is in comparison [39, 40] marginally less design centric. However, to improve detection resolution, it may use the know-how of direct sensor-actuator relationships, available from the testbed design document or derived from the system architecture by a subject expert, and PLC imposed non nn-actuation (see Section VI-D) logic available in the PLC control logic sheet. Note, we do not break it down to process stages or look for invariants across sub-systems to find association rules. Instead, boundaries are determined per sensor and relationships are generally constrained to nn-actuator(s). We assume that a suitable approach was used to set the PLC min-max boundaries discussed in Section II-A. These hard boundaries are essential for correct system operations even in the absence of anomalies. Recent developments in using large language models (LLMs) for anomaly detection is extensively reviewed in Su et al [41]. Unlike GiBy, none of the log analysis detectors surveyed there positioned themselves as real-time detectors. A generalized note they made on computational efficiency in LLM deployment for forecasting and anomaly detection was its sheer scale and complexity of those models. It demanded substantial computational resources, likely to limit its scalability and usability for real-time applications. Also of concern was the sustainability and energy efficiency of LLMs with future trends likely to persist on environmentally friendly models through practices such as optimized algorithms and energy-efficient hardware. In this aspect, unlike LLMs and some of the deep learning AI models discussed; GiBy showed millisecond detection speeds on an everyday used microprocessor demonstrating that the algorithm is energy efficient.
VIII Conclusions and Future Work

### v3

@@ 6.1 Detector limitations
We mapped a smaller state-space of actuator states to the sensor reading. As a result, we forsook the determination of a larger number of tighter individual bounds that takes exponential time for the convenience of a practically smaller solution seek space. However, our sensor measurements are taken directly from the normal operations dataset. As a result, the overall determined [LB, UB] bounds remained unchanged, and hence does not change the overall safety limits determined. For example, consider eight dependent actuators influencing a sensor reading. When the actuation sequence for each actuator is on, off, then . When and number of actuators is 8, there are a total of actuation states. This requires determination of bounds. Additionally, let us assume that only 3 out of the 8 dependent actuators are nearest neighbours, w.r.t. sensor . Using GiBy, we would have only considered nearest neighbour () actuation states, and determined . As a result, there was no change in the overall state sensor bounds (safety limits) determined, although some detection resolution was lost by considering only the nearest neighbours.
@@
Another limitation expected from the anomaly detector is a sensitivity to out of bound measurements, resulting in warnings being issued when no attack is present. This limitation could occur even in the system normal (initial) operating phase. The normal dataset is expected to cover its everyday boundaries of operation. However, in some cases the normal dataset may not cover all boundaries of normal operations, and these normal boundaries may be surpassed during future plant operations. Any logged warning observed, that is deemed as normal, will have to be reviewed and trained on, to avoid those warnings from re-appearing. However, this will stabilize once sufficient normal (training) data indicative of its range of operations is available. This is a training dataset limitation that affects our model. A final limitation is that it may not detect all invalid actuation states. It is limited to the resolution of the actuator states seen for the nn-actuators. Caution must be exercised in systems where actuators that are not nearest neighbors; whose actuation results in unsafe operations — may not be detected by GiBy using only nn-actuators. The training will need to include those other actuators that are in relationship with the sensor.
6.2 Detection in an early prototype system design
@@
6.3 GiBy Detection capabilities
@@
In the threat model (Section 2.4), the attacks are classified into single-stage single/multiple component attacks and multi-stage single/multiple component attacks. The component considered is a sensor or an actuator. The goal of this classification is to build a quantifiable measure (score) on the quality of the detector, w.r.t. its ability to detect these types of attacks. Without loss of generality, the classification may be merged into single and multiple sensor/actuator attack(s) on a system.
@@
Our core detector (see Section 3) can detect single sensor or actuator manipulations, when it breaks the safety bounds of the relation between sensor and its actuation state. When both sensor and corresponding actuator(s) are spoofed, it is no longer able to individually detect that attack, but it depends on whether any of the sensor readings downstream222Downstream bounds are also captured by GiBy, since it is computed for every sensor with nn-neighbour actuator(s)., as a result, went out of bounds. For example, in Fig. 4, consider an attack where P101 is spoofed to ON and FIT201 reading to , despite the pump being OFF and no water flowing. Assume the controller logic is to turn on the first HCl dosing pump (P201), when the FIT201 flow rate is between and to OFF otherwise. The attack causes the dosing pump to release HCl even when no water is pumped333The success of this attack depends on whether the attacker can compromise the controller PLC and manipulate its measurements.. However, due to pH sensor AIT201, this attack is detected. Further, if we assume that AIT201 is also spoofed to show normal range of measurements — a redundant downstream sensor AIT501 incorporated in is able to detect the abnormality in pH. However, if AIT501 is also spoofed, it may not be able to detect the attack from the measurements made. As a side effect, changing the pH might possibly change its conductivity sensor reading, permitting the attack to be detected. It would be difficult to conclude without further experimentation. In some scenarios, it may be sufficient to detect the anomaly at a later stage, and in others, it requires to be detected in the same stage. For example, if pump P101 is running above its safe rpm, it might cause the adjacent (following) pipe to burst, affecting the plant and personnel. Where such a concern exists, the design has to include a pressure sensor. When there is a risk the pressure sensor reading is spoofed, a physical pressure safety (release) valve will have to be incorporated into the design [25]. An anomaly detector’s ability to detect single/multiple attacks is highly dependent on intrinsic design relationships and system dependencies. This tight coupling makes generalized interpretations on detection capability (by only considering the detection model) less meaningful.
6.4 Non-neighbor PLC imposed actuation
@@
The stage in Fig. 4 involves chemical dosing. Dosing tanks are required to be in the order NaCl, HCl and NaOCl (see Section 6.2 for reason). AIT201 in the SWaT dataset measures conductivity. According to the PLC control logic sheet, the conductivity sensor AIT503 (downstream sensor in ) read by the PLC instructs P201 to be turned off, when this sensor value hits a preset threshold. Turning off the salt pump, in-turn regulates the conductivity sensor reading at AIT503. Hence, P201 is additionally considered as a nn-actuator (though not a neighbour) for the sensor AIT503; to capture this relationship. It is added to train the sensor (using GiBy). In general, the PLC logic sheet is examined to include such ‘additional’ nn-actuators, when there is a sensor dependency. Not including this relationship does not change the overall detection bounds captured in training. However, it means that some detection resolution is lost because we are no longer directly capturing the relationship between AIT503 and pump P201.
6.5 Explainablity
@@
The more sophisticated AI/ML models using neural network and deep learning are inherently black box models. They are used to detect anomalies but require additional support from eXplaianable AI (XAI) models to explain the decision made by the AI detector. However, both the ML and XAI models have individual accuracy limitations and this is compounded when used in combination. It also takes longer to train and test these models (see Section 5). On the contrary, explainability with the proposed GiBy detector is straightforward. Note that we employed a switchboard in Section 3.3. Traceability is offered via the one-to-one and onto map, using the switchboard — also in the inverse direction, from sensor reading (under testing) to its [LB, UB] bounds (determined in the training phase). Our training and testing are carried out per sensor, for all concerned sensors and nn-actuator(s). When an anomaly is detected, this immediately tells us what sensor readings are causing anomalies. Since the sensor is known, we can look at its current and previous actuation states, from the state log. GiBy raises an anomaly warning when i.) the sensor reading (under testing) for the nn-actuator state is out of bound. Our detector is able to provide an explanation remark, whether it was because the sensor reading was or . ii.) when the normal training dataset did not have the nn-state actuator state for the sensor, in it. This may happen for two reasons a.) the actuation state is invalid or b.) that state is yet to be encountered in the plants’ normal operation and hence not yet seen in the normal training dataset. Our detector system is designed to offer a level of explainability that is expected to equip the operator to deal with most contingencies. The capability of our system to provide the correct sensor location and response to faulty components/attacks subsequently reduces the plant downtime.
6.6 Operator Workload Reduction
We revisit the definitions of precision, recall, accuracy, and F1-score in Equations (6)-(9), and discuss what it means for the operator.
In our case, GiBy is an one-class classifier and it considers all sensor readings within the determined sensor boundaries as safe. For the purpose of reducing the operator workload, we define (for this section) any sensor reading within the safety boundary to be non-anomalous. This implies that even if a sensor reading had changed but its resulting value was within the safety limits; it is declared a safe and non-anomalous operation. A false positive is recorded when the model incorrectly predicts a positive outcome (within safety boundary) as an outcome that is negative (outside safety boundary). For example, a false positive occurs when a safe sensor reading value is predicted as unsafe (anomalous). Training is carried out on a normal dataset to determine safety boundaries. The safety boundaries expand when further training data is made available; and eventually the rate of expansion stabilizes — once it has seen most of the normal operations boundaries. A false positive limit may be chosen by the operator and GiBy may be used once it meets this requirement. This allows an operator to manage the number of FPs seen.
A false negative is recorded when the model incorrectly predicts a negative outcome (outside safety boundary) as an outcome that is positive (within safety boundary). In GiBy, there are no FNs because as it continues to train on normal data, the safety boundary expands outwards from a point.
Since GiBy has no FN and its FP is determined by the length and breadth of the boundaries seen in normal training data — we argue that GiBy’s true positive and true negative are a measure of its detection capabilities. Note that in Equation (6), as 0, 1. In Equation (7), as 0, 1. For GiBy, as more training data is available, its 0 and 1. Since there are no FNs in GiBy, its . This implies, an extensively trained GiBy that has seen most boundaries will have and 1, resulting in a F1-score close to 1; w.r.t. the TP, FP and FN the operator sees.
When FP (as is the case for it to be practically useful for the operator), its precision is also dominated by the TP. The operator is now tasked with resolving actual anomalies and the occasional FP. The FP frequency of occurrence is loosely bounded by the false positive limit (threshold). This is a configurable parameter set by the implementer during setup. However, it is required to meet the average FP frequency threshold set for usability.
6.6.1 GiBy Re-defined Accuracy
When we used the re-definition for accuracy shown in Equation (8) — it additionally considers any anomalous (changed) reading within the safe operation limits as not anomalous, to reduce the operator workload. Further, it is straight-forward to report any state not seen in training as anomalous/not seen. Now, with reference to TABLE IV, attacks 13, 14, 17, 24, 27, and 29, will be considered as safe and non-anomalous. Only attack 3 is undetectable by GiBy. The re-defined accuracy is 43/44 = 0.9772 or 97.72%. We maintain that the detection accuracy of 78.5% (based on a conventional444For example, consider an email spam filter model, that correctly classifies 90 out of 100 emails. Its detection accuracy is 90/100=0.9 = 90%. Even for a single email with 1000 detected spam words, it is counted as one spam email, unless a contextual reasoning for re-defining accuracy is provided. definition [26]) reported in Section 5.0.1 is high. This is because GiBy correctly detected all anomalies outside its normal (seen) trained bounds.
6.7 Anomaly Detection Woes: Improving detection accuracy by using alternate definitions and experiments
In this section we highlight certain techniques used to increase detection accuracy by changing the definition for accuracy seen in Equation (8). We also discuss serious implications from inflated reporting of detection accuracy and how it affects the works that define and report accuracy in the conventional manner; especially when they are expected to be compared with “state of the art” work claiming high accuracies in the 95-99.99% range. The purpose of this discussion is to highlight the different definitions for the seemingly same reference to the word “detection accuracy”, and why some works in existing literature are not comparable with GiBy’s detection accuracy. The interpretation of what accuracy measures and how it is reported is inconsistently defined and calculated; making GiBy detection accuracy comparisons incompatible with some works in existing literature. The work in Inoue et al. [27] reports using SWaT dataset for their DNN model, a precision of 0.98295. For One-class SVM, their reported precision is 0.92500. It reads – “It should be noted that false positive and true positive rates, which underlie these statistics, are counted over log entries for DNN whereas they are counted over windows for SVM”. In GiBy, the accuracy is determined by Equation (8) (also see Section 5.0.1 and Section 6.6.1). This is incompatible with the definition of log entries and windows used. The work in Xu et al. [28] reports SWaT detection accuracy for different models ranging from 95.97% to 99.95% (refer to their attack classification performance on the SWaT Dataset table). It reads — “After downsampling, the SWaT dataset contains 47044 samples, including 26877 normal samples”. However, for GiBy — we use the full range of values in the training and testing dataset discussed in Section 2.3. Further, there are a total of 41 attacks on SWaT shown in TABLE IV. Even if 40 out of 41 attacks were detected, the accuracy would be . To arrive at an accuracy of 0.99, the division has to be no less than 99/100 = 0.99. In Xu et al. [28], they report accuracies up to 0.9995, which is a mathematical impossibility; for how accuracy is defined for GiBy (see Section 5.0.1 and Section 6.6.1). This clearly indicates a different definition was used for their detection accuracy; making any comparisons with GiBy incompatible.
6.8 Usability in other sectors
@@
Anomaly detection and mitigation in some sectors (such as power), in certain scenarios, require an actuation within 10-15 milliseconds from event occurrence, to ensure safety. The GiBy algorithm presented in Section 3 is universally transferable and GiBy-core testing may be parallelized. This is because it produces a series of independent [LB, UB] boundaries. It may be pre-programmed into the control logic of an Intelligent Electronic Device (IED), typically used in a power substation. It may be used to trip the circuit breaker in near real-time (when sensor readings are out of safety bounds), or to inform an operator regarding the anomalous behaviour. The computational cost of GiBy-core tests (see Section 3.5) are to see whether the sensor value is in between two bounds [LB, UB]; taking up to two comparisons. The main costs of the extended algorithm tests (see Section 4) are the binary search and sliding window multiplications in each time step. Its time complexity may be reduced to that of binary search on the pre-sorted linearized group (that takes logarithmic time), and a maximum of one multiplication and division: for each time step. Instead of repeatedly multiplying every sensor derived value in the sliding window, the newest time step sensor derived value may be multiplied into the existing product, and the sensor derived value at the tail of the window is divided out. Due to the straightforward mappings and simple boundary checks in GiBy, it might also find (outlier detection) applications in resource constrained devices, edge computing devices and legacy systems with lower computational power.
7 Related Work
@@
The focus of this work is on process anomaly detection [29, 30]. Several approaches are available for design-centric and data-centric anomaly detection. In a design-centric approach, for example, Adepu and Mathur [31] captured the design using state condition graphs. It was used to capture the conditions on actuation w.r.t. sensor readings, expressed using Boolean conditions, and represented in a graphical format. The work in Yoong et al. [32] used axiomatic design methodology for systems, where it iteratively decomposes a CPS design to sets of dependent components and transformed into invariants. The work in Raman and Mathur [33] has a detection rate of over 90% against actuators and above 75% against sensors, excluding certain stealth attacks. In GiBy, out of bound sensor values are considered as anomalies; irrespective of whether they were indirectly caused by changes in actuator state or directly by altering the sensor value. In Merwa et al. [34], an association rule mining technique was used to generate attack and invariant rules. In a data-centric approach, for example, there exist solutions to distinguish outliers, such as ECOD in Li et al. [10]. It computed a univariate empirical cumulative distribution function for each dimension separately. To measure the chance of a data point being anomalous, they computed its tail probability across all dimensions. Another outlier detector called deep Support Vector Data Description (Deep-SVDD) proposed in Ruff et al. [11], trained a neural network while minimizing the volume of a hypersphere that enclosed the network representations of the data. In WaXAI, Mathuros et al. [12] employed ECOD and Deep-SVDD in the context of SWaT anomaly detection. In addition, they also employed XAI [14] models, namely, kernel SHAP [35], LIME [36], ALE [37] and IG [38], to provide explanation for the anomaly detected. Further, a 20% improvement in attribution of attack root cause over SHAP was achieved by using a Factorization Machines based approach by Avdalovic et al. [39]. In this work we provided attribution (explainability) linked directly to the sensor. Explainability was provided for each sensor when out of bounds, to top up the recommendations in Fung et al [40]. GiBy showed millisecond detection speeds on an everyday used microprocessor demonstrating that the algorithm is energy efficient.
8 Conclusions and Future Work